Consider the following data:

x 5 6 7 8 9
P(X=x)P(X=x) 0.3 0.1 0.1 0.3 0.2

Copy Data

Step 2 of 5 :

Find the variance. Round your answer to one decimal place.
Answer #1

I var(x) = E(XP) - (E(X))2 -- formula E(X) = Expca) = 5x0.3 +6x0,1 +7x0.1 +8x0.3 +9x12 - 1570.6 +0.7+254 +1.8 E(x² = ž x² p(x

Answer = 2.4